If the ratio of 3x - 7y + 4 to 4x + 3y + 12 is 1/3 then the ratio of x to y is
A –18/5 b 5/18 c 2 d 24/5 e none of these

Answer by anantha(86)

sol
the ratio of 3x-7y+4 to 4x+3y+12 is 1/3 means
(3x-7y+4)/(4x+3y+12)=1/3
on cross multiplication
3*(3x-7y+4)=1*(4x+3y+12)
9x-21y+12=4x+3y+12 for isolating x we subtract both sides by 4x,3y,12
9x-21y+12-4x-3y-12=4x+3y+12-4x-3y-12
5x-24y=0
again add 24y on both sides
5x-24y+24y=0+24y
5x=24y
dividing both sides by 5
5x/5=24y/5
x=24y/5
again dividing by y
x/y=24y/5y
x/y=24/5
the answer is d)